# Steiner Traveling Salesman Problem with Time Windows and Pickup–Delivery: integrating classical and quantum optimization

### v2

@@ 6.4.2 Performance Analysis with a quantum solver
The results reported in Table 14 highlight the impact of excluding pickup operations on the behavior of the ABF model, both with and without the AFGR mechanism. For small-size instances ($V=4$), the performance is already strong even without AFGR: the average optimality gap is close to the reference value ($GAP=4\%$) with 100% of instances solved. Nevertheless, the inclusion of AFGR further improves the results by reducing the average gap to 2% and ensuring greater stability, as reflected in the lower standard deviation. In this case, the slight increase in average computational time is negligible and fully justified by the gain in robustness. When the problem size increases ($V=5$), the model becomes less stable: the gap rises to 28% without AFGR, with high variability across runs (Std=167,55). The introduction of AFGR reduces the average gap to 23% and increases solution consistency, although variability is not entirely eliminated. Moreover, the average computational time decreases significantly (from 54,48s to 39,15s). For $V=6$ and $V=7$, the problem becomes more challenging. Without AFGR, the gap reaches 51–60%, with only 60% and 30% of instances solved, respectively. The integration of AFGR does not reduce the gap, which remains high, but it considerably improves the percentage of solved instances (80% for $V=6$ and 50% for $V=7$) and leads to more stable outcomes, as shown by the reduction in standard deviation. This indicates that in more complex scenarios AFGR plays a crucial role primarily in ensuring feasibility and reducing variability rather than directly improving the objective function value. For larger instances ($V\geq 8$), the difficulty further increases. Without AFGR, the algorithm fails to produce any feasible solution (0% solved). With AFGR, non-optimal but consistent solutions are obtained: for $V=8$ the average gap is 57% with 40% of instances solved, while for $V=9$ the gap is 47% with 30% solved. These results confirm the essential role of AFGR as a recovery mechanism for feasibility in highly complex cases. Finally, for $V=10$, neither configuration is able to consistently generate valid solutions. The problem becomes intractable within the given time limits, suggesting that additional acceleration or decomposition strategies would be necessary to deal with instances of this scale.
Overall, the analysis shows that AFGR becomes increasingly beneficial as
@@ 7 Conclusions and Further Work
This work introduces the STSP-TWPD, a novel and practically relevant extension of the classical STSP. By integrating time windows, pickup and delivery operations and vehicle capacity constraints within a Steiner graph structure, the STSP-TWPD captures the multifaceted challenges faced in real-world logistics applications such as last-mile delivery, reverse logistics, and time-sensitive distribution. To address the inherent computational complexity of this NP-hard problem, we develop two original mathematical formulations: an ABF that extends traditional STSP models by explicitly incorporating temporal and logistical dimensions, and a NBF, a novel modeling We also introduce formal mathematical formulations for the STSP-PD and the STSP-TW, which, to the best of our knowledge, have not previously appeared in the literature. Furthermore, we propose a heuristic reduction method, the AFGR, a preprocessing technique that reduces problem dimensionality by eliminating redundant arcs, while preserving feasibility, significantly enhancing scalability and solver performance. We solved the proposed models to optimality using Gurobi and additionally assessed the feasibility and compatibility of the STSP-TWPD formulations within a hybrid quantum–classical computing framework. By implementing our models on D-Wave’s LeapCQMHybrid solver, we The experimental results confirm the effectiveness of the AFGR method in reducing model size by more than 40% on average, improving solution feasibility and stability across a wide range of configurations.
@@
Data will be made available on request.

### v3

@@ 6.4.2 Performance Analysis with a quantum solver
The results reported in E.3 (Table 48) highlight the impact of excluding pickup operations on the behavior of the ABF model, both with and without the AFGR mechanism. For small-size instances ($V=4$), the performance is already strong even without AFGR: the average optimality gap is close to the reference value ($GAP=4\%$) with 100% of instances solved. Nevertheless, the inclusion of AFGR further improves the results by reducing the average gap to 2% and ensuring greater stability, as reflected in the lower standard deviation. In this case, the slight increase in average computational time is negligible and fully justified by the gain in robustness. When the problem size increases ($V=5$), the model becomes less stable: the gap rises to 28% without AFGR, with high variability across runs (Std=167,55). The introduction of AFGR reduces the average gap to 23% and increases solution consistency, although variability is not entirely eliminated. Moreover, the average computational time decreases significantly (from 54,48s to 39,15s). For $V=6$ and $V=7$, the problem becomes more challenging. Without AFGR, the gap reaches 51–60%, with only 60% and 30% of instances solved, respectively. The integration of AFGR does not reduce the gap, which remains high, but it considerably improves the percentage of solved instances (80% for $V=6$ and 50% for $V=7$) and leads to more stable outcomes, as shown by the reduction in standard deviation. This indicates that in more complex scenarios AFGR plays a crucial role primarily in ensuring feasibility and reducing variability rather than directly improving the objective function value. For larger instances ($V\geq 8$), the difficulty further increases. Without AFGR, the algorithm fails to produce any feasible solution (0% solved). With AFGR, non-optimal but consistent solutions are obtained: for $V=8$ the average gap is 57% with 40% of instances solved, while for $V=9$ the gap is 47% with 30% solved. These results confirm the essential role of AFGR as a recovery mechanism for feasibility in highly complex cases. Finally, for $V=10$, neither configuration is able to consistently generate valid solutions. The problem becomes intractable within the given time limits, suggesting that additional acceleration or decomposition strategies would be necessary to deal with instances of this scale.
Overall, the analysis shows that AFGR becomes increasingly beneficial as instance complexity increases. For the smallest cases, it yields only marginal improvements in solution quality, whereas for more challenging configurations it plays a crucial role in obtaining feasible and stable solutions. AFGR also consistently reduces variability, even when the optimality gap remains significant, thereby enhancing the robustness of the solving process. Nevertheless, the limitations of current quantum technology become evident as the problem size increases: beyond $V=8$, even with AFGR, the solutions deviate substantially from the optimum and the percentage of solved instances decreases. This confirms the role of AFGR as an effective support mechanism for maintaining feasibility and stability, while acknowledging that overall scalability remains constrained by the capabilities of present hybrid quantum–classical solvers.
@@ 7 Conclusions and Further Work
This work introduces the STSP-TWPD, a novel and practically relevant extension of the classical STSP. By integrating time windows, pickup and delivery operations and vehicle capacity constraints within a Steiner graph structure, the STSP-TWPD captures the multifaceted challenges faced in real-world logistics applications such as last-mile delivery, reverse logistics, and time-sensitive distribution. To address the inherent computational complexity of this NP-hard problem, we develop two original mathematical formulations: an ABF that extends traditional STSP models by explicitly incorporating temporal and logistical dimensions, and a NBF, a novel modeling representation for the STSP. We also introduce formal mathematical formulations for the STSP-PD and the STSP-TW, which, to the best of our knowledge, have not previously appeared in the literature. Furthermore, we propose a reduction method, the AFGR, a preprocessing technique that reduces problem dimensionality by eliminating redundant arcs, while preserving feasibility, significantly enhancing scalability and solver performance. We generated an extensive set of benchmark instance classes specifically designed to systematically test the proposed models and solution methodologies under varying levels of complexity. We solved the proposed models using classical solver and additionally assessed the feasibility and compatibility of the STSP-TWPD formulations within a hybrid quantum–classical computing framework. By implementing our models on D-Wave’s LeapCQMHybrid solver, we provide a proof of concept demonstrating that complex routing problems such as the STSP-TWPD can already be represented and solved, at least for small instances, using emerging quantum optimization paradigms. The experimental results demonstrate that the AFGR method can reduce the model size by more than 70%, while consistently achieving objective values that are equal to or even better than those obtained without AFGR.
@@
The data used in this study are available in the repository provided in this work. The benchmark datasets employed are included and referenced within the repository.